[Brahms-l] Shift report 20040226 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Thu Feb 26 2004 - 23:55:05 EST
------------------------------------------------------------
Date: Thu Feb 26 23:55:05 2004, Shift: 16:00-24:00
Supervisor: DO, with: 
Log book#15 pages 67 - 67

Shift Summary:
------------------------------------------------------------
1. End of store at beginning of shift (16.30), new store at 
   18.45 after RHIC access requested by STAR. Lasted until 
   23h15. -> 5 hours of beam

2. Data: started MRS 40 B 1/1, goal is 2M tr3
   MRS 40 B1/1 : 478k tr3 (runs 10616 -> 10622)
   FS   8 A1/2 : 217k tr2 (runs 10616 -> 10623)

   Took a short MRS 0 field run at end of store (83k tr3)

   Pedestal runs: 10617 & 10624 (TPC)  and 
                  10619 & 10625 (non TPC)

3. None. Flemming set the FS trigger back to its
   original setting (7 slats)

4. This was my last shift ;D yeeeees!



 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 26 23:55:16 2004

This archive was generated by hypermail 2.1.8 : Thu Feb 26 2004 - 23:55:36 EST